OpenStudy (anonymous):

What is the equivalent resistance of a parallel circuit with a 5.0 kΩ resistor, a 2.5 kΩ resistor and a 10.0 kΩ resistor? A. 1.4 kΩ B. 0.10 kΩ C. 0.83 kΩ D. 10 kΩ E. 18 kΩ

OpenStudy (michele_laino):

the equivalent resistance is given by the subsequent formula: \[\Large {R_{EQ}} = \frac{1}{{\frac{1}{{2.5}} + \frac{1}{5} + \frac{1}{{10}}}} = ...K\Omega \]

OpenStudy (anonymous):

ok! so 1.43? so choice A is our solution?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay!! thanks!:)

OpenStudy (michele_laino):

:)